# Otter: Generating Tests from Issues to Validate SWE Patches

### v1

@@ 5.2 Effectiveness of Otter, Otter++, and the Baseline
Table 1 shows that GPT-4o-based Otter and Otter++ perform well on test generation, creating fail-to-pass tests for 31.4% and 37% of the instances, respectively, whereas the baseline produced such tests for 18.7% of the instances. The improvements are also reflected in , a metric that blends fail-to-pass rate with coverage (Ahmed et al., 2024). We observe similar performance improvements with Mistral-large. The pass@5 rate (where one of the five tests is fail-to-pass) for Otter++ is 44% for GPT-4o and 37% for Mistral-large.
@@
tddScore | ||||
| Mistral-Large | Zero-shot | 57 | 12.7 | 11.8 | ||||
| Otter | 121 | 26.9 | 25.1 | |||||
| Otter++ | 144 | 32.1 | 28.6 | |||||
| GPT-4o | Zero-shot | 99 | 18.7 | 17.2 | ||||
| Otter | 141 | 31.4 | 29.4 | |||||
| Otter++ | 166 | 37.0 | 32.4 |
@@ 5.3 Ablation Study
| Model | Approach | # of fail-to-pass | tddScore |
| Mistral-large | Otter (T1) | 121 | 25.7 | NA | ||
| Otter – Action Planning* (T2) | 96 | 20.2 | -21.4 | |||
| Otter – Focal Localization* (T3) | 96 | 20.2 | -21.4 | |||
| Otter – Test Localization* (T4) | 77 | 16.5 | -35.8 | |||
| Other - Focal & Test Localization* (T5) | 81 | 17.2 | -33.1 | |||
| Otter – Fixing Import | 117 | 24.6 | -4.3 | |||
| Otter – Imports at Generation | 114 | 23.7 | -7.8 | |||
| GPT-4o | Otter (T1) | 141 | 29.4 | NA | ||
| Otter – Action Planning* (T2) | 110 | 23.6 | -19.7 | |||
| Otter – Focal Localization* (T3) | 115 | 25.2 | -14.3 | |||
| Otter – Test Localization* (T4) | 107 | 24.2 | -17.7 | |||
| Other - Focal & Test Localization* (T5) | 110 | 24.2 | -17.7 | |||
| Otter – Fixing Import | 128 | 26.7 | -9.2 | |||
| Otter – Imports at Generation | 130 | 26.7 | -9.2 | |||
| * is not followed by action planning | ||||||
@@
Mündler et al. (2024) proposed a set of approaches for generating fail-to-pass tests. We ran Otter and Otter++ on their dataset to study how the approaches compare. They also evaluated zero-shot approaches, which differ from our zero-shot baseline. Instead of generating a complete function, all of their approaches (including the zero-shot ones) instruct the model to generate a specific form of “diff”. Two of their approaches use a golden patch in the prompt. Mündler et al.’s SWE-agent and SWE-agent+ approaches are derived from SWE-Agent, which was originally designed for generating golden code patches (Yang et al., 2024). Table 3 shows the results. Otter and Otter++ perform better than their best-performing approach, generating 70 (25.4%) and 80 (29.0%) fail-to-pass tests compared to 53 (19.2%) fail-to-pass tests generated by SWE-agent+.
@@
| ZeroShot | 16 | 5.8 | ||
| ZeroShotPlus* | 28 | 10.1 | ||
| LIBRO* | 42 | 15.2 | ||
| AutoCodeRover | 25 | 9.1 | ||
| SWE-Agent | 46 | 16.7 | ||
| SWE-Agent+ | 53 | 19.2 | ||
@@
| * follows “write first, test later” approach | ||||
@@ 5.5 Cost Effectiveness of Otter
Table 4 presents the cost for invoking the GPT-4o model to process 449 instances with Otter. Each sample requires an average of $0.06. The “reflect and improve the plan” step can make 1–5 LLM calls. Figure 2 shows that, for more than 80% of the instances, both GPT-4o and Mistral are satisfied within two calls. Therefore, the total calls vary from 7–8 for most instances. The cost is very low because we do not accumulate context from prior calls in subsequent calls. We do not discuss the cost for Mistral-large because the model was hosted locally. Otter++ makes additional four calls and reuses the output from localizers. The total cost for Otter++ (which includes Otter) is $0.09 per instance.
@@ 5.6 Characteristics of the Generated Tests
Figure 3 presents the overlap among instances for which fail-to-pass tests could be generated by different prompts using the GPT-4 model. Overall, each of the prompts is successful on some instances on which the other prompts fail, with T1 (Otter with GPT-4o) achieving the most success in this respect—producing fail-to-pass tests for 24 instances on which none of the other prompts succeeded in generating such tests. Thus, combining the results from the different prompts increases the fail-to-pass rate to 44% (38% for Mistral) at pass@5. This indicates the benefit of incorporating different prompts for test generation.
@@ 5.7 Test Generation and SWE Agents
The tests generated by our approach can be used for filtering bad code patches and increasing the precision of solutions proposed by different systems from the SWE-bench Verified leaderboard. We take the top 22 systems from the leaderboard and run the five tests generated by Otter++. We filter out a code patch if all the tests fail on it. Figure 4 shows that we have achieved a precision of 65% to 92% while maintaining a decent recall of 30%-41%, except for one system where the precision increased by 22% to 167%. Note that Mündler et al. (Mündler et al., 2024) achieved 47.8% precision at 20% recall on SWE-Agent. Using tests generated by Otter++ achieved much higher precision while maintaining greater recall with 22 systems.
@@ 5.8 Effect of Data Contamination
We conducted another experiment following the approach of Schäfer et al. (2023). We take the generated test from Otter and compute the similarity score with the most similar test from the repository. We compute the score between the generated test and all existing tests as follows: , where is the set of test functions and is the generated test. Figure 6 shows the similarity scores for new tests. For 90% of the instances, the similarity score is less than 0.6. We share some samples with more than 0.5 similarity in Table 8 (in the appendix) to give the reader some idea. From our observation, even at similarity score of 0.7, the tests are significantly different. As expected, for modified tests, the similarity is higher. However, we did not observe any difference between fail-to-pass tests and other tests (see Appendix for figures). Therefore, we believe that the model is not simply generating memorized tests.
@@ 7 Conclusion
The primary contribution of this paper is Otter, a system for generating tests from issue descriptions before issue resolution. Otter outperforms the prior state of the art in fail-to-pass tests generated while also costing less. Furthermore, this paper demonstrates that generated tests can improve patches generated by SWE agents, helping them reach a precision of between 65% and 92%.
@@
This paper presents work whose goal is to advance the field of software testing and program repair. Developers spend a significant amount of time resolving bugs and testing them. We believe this work will significantly improve the developers’ experience in their day-to-day life.
@@ Appendix A Different attributes of the TDD-Bench-Verified Instances & Sample Test Patch
| *File counts are based on the main branches of the project (cloned on October 29, 2024). | |||||||||
@@ Appendix B Action Counts in Self-Reflective Action Planner Phase
Appendix C Data Contamination
@@
Appendix D Prompts for Otter

### v2

@@ 6.2 Effectiveness of Otter, Otter++, and the Baseline
Table 1 shows that GPT-4o-based Otter and Otter++ perform well on test generation, creating fail-to-pass tests for 31.4% and 37% of the instances, respectively, whereas the baseline produced such tests for 18.7% of the instances. The improvements are also reflected in . We observe similar performance improvements with Mistral-large. The pass@5 rate (where one of the five tests is fail-to-pass) for Otter++ is 44% for GPT-4o and 37% for Mistral-large.
@@
tddScore | Coverage | ||||
Mistral-Large | Zero-shot | 57 | 12.7 | 11.8 | 60.6 | ||||
Otter | 121 | 26.9 | 25.1 | 70.5 | |||||
Otter++ | 144 | 32.1 | 28.6 | 70.4 | |||||
GPT4o | Zero-shot | 84 | 18.7 | 17.2 | 60.0 | ||||
Otter | 141 | 31.4 | 29.4 | 70.6 | |||||
Otter++ | 166 | 37.0 | 32.4 | 71.5 |
@@ 6.3 Ablation Study
Model | Component | Approach |
tddScore |
Mistral- large | - | Otter (T1) | 121 | 25.7 | - | ||||
Action Planner | without Action Planning (complete)* (T2) | 96 | 20.2 | -21.4 | |||||
without Plan Refinement (Just 1 attempt) | 115 | 23.8 | -7.4 | ||||||
without Action Validation | 107 | 22.1 | -14.0 | ||||||
Localizers | without Focal Localization* (T3) | 96 | 20.2 | -21.4 | |||||
without Test Localization* (T4) | 77 | 16.5 | -35.8 | ||||||
without Focal & Test Localization* (T5) | 81 | 17.2 | -33.1 | ||||||
Test Generator | without Fixing Import | 117 | 24.6 | -4.3 | |||||
without Imports at Generation | 114 | 23.7 | -7.8 | ||||||
GPT-4o | - | Otter (T1) | 141 | 29.4 | - | ||||
Action Planner | without Action Planning (complete)* (T2) | 110 | 23.6 | -19.7 | |||||
without Plan Refinement (Just 1 attempt) | 130 | 27.5 | -6.5 | ||||||
without Action Validation | 120 | 25.7 | -12.6 | ||||||
Localizers | without Focal Localization* (T3) | 115 | 25.2 | -14.3 | |||||
without Test Localization* (T4) | 107 | 24.2 | -17.7 | ||||||
without Focal & Test Localization* (T5) | 110 | 24.2 | -17.7 | ||||||
Test Generator | without Fixing Import | 128 | 26.7 | -9.2 | |||||
without Imports at Generation | 130 | 26.7 | -9.2 | ||||||
* is not followed by action planning |
@@
Mündler et al. (2024) proposed a set of approaches for generating fail-to-pass tests. We ran Otter and Otter++ on their dataset to study how the approaches compare. They also evaluated zero-shot approaches, which differ from our zero-shot baseline. All of their approaches (including the zero-shot ones) instruct the model to generate a novel code diff format introduced by their paper. Two of their approaches use a proposed patch in the prompt. Mündler et al.’s SWE-agent and SWE-agent+ approaches are derived from SWE-Agent, which was originally designed for generating code patches (Yang et al., 2024). Table 3 shows the results. Otter and Otter++ perform better than their best-performing approach, generating 70 (25.4%) and 80 (29.0%) fail-to-pass tests compared to 53 (19.2%) fail-to-pass tests generated by SWE-agent+.
Otter++ uses execution logs on the current code base which give it an advantage in the final selection stage. The feedback works as a contributing factor to the superior performance of Otter++. Note that Otter does not use any execution logs, yet it still performs significantly better than LIBRO (25.4% vs. 15.2% in Table 3). The novel code diff format proposed by Mündler et al. has similarities to our approach. However, their format requires the model to perform additional tasks such as writing the file name, change type, and line number in response to one LLM call. Additionally, to instruct the model to generate a specific format, the authors had to include an example that is not relevant to the issue itself. In our generation step, the model only needs to generate the test (and prior function name for positioning new tests).
@@
ZeroShot (GPT-4) | 16 | 5.8 | ||
ZeroShotPlus* (GPT-4) | 28 | 10.1 | ||
LIBRO* (GPT-4) | 42 | 15.2 | ||
AutoCodeRover (GPT-4) | 25 | 9.1 | ||
SWE-Agent (GPT-4) | 46 | 16.7 | ||
SWE-Agent+ (GPT-4) | 53 | 19.2 | ||
@@
* uses “proposed patch” while generating tests |
@@ 6.5 Cost Effectiveness of Otter
Table 4 presents the cost for invoking the GPT-4o model to process 449 instances with Otter. Each sample requires an average of $0.06. The “reflect and improve plan” step can make 1–5 LLM calls. Figure 4 shows that, for more than 80% of the instances, both GPT-4o and Mistral are satisfied within two calls. Therefore, the total calls vary from 7–8 for most instances. The cost is very low because we do not accumulate context from prior calls in subsequent calls. We do not discuss the cost for Mistral-large because the model was hosted locally. Otter++ makes four additional calls and reuses the output from localizers. The total cost for Otter++ (which includes Otter) is $0.09 per instance.
@@ 6.6 Characteristics of the Generated Tests
Figure 5 presents the overlap among instances for which fail-to-pass tests could be generated by different prompts using the GPT-4 model. Overall, each of the prompts is successful on some instances on which the other prompts fail, with T1 (Otter with GPT-4o) achieving the most success in this respect—producing fail-to-pass tests for 24 instances on which none of the other prompts succeeded in generating such tests. Thus, combining the results from the different prompts increases the fail-to-pass rate to 44% (38% for Mistral) at pass@5. This indicates the potential of incorporating different prompts for test generation.
Detailed analysis of the tests.
Table 6 presents the analysis of the Otter-generated tests using GPT-4o model from different perspectives. We see that the fail-to-pass rate is higher when patching an existing test than when adding a new test, which is expected because the model can have better context from the existing test. Writing a new test is inherently more difficult than modifying an existing test. It is expected that the test will fail on the old codebase. However, we found 69 samples in Otter where the test passed on the old codebase. Also, our analysis shows that tests with assertion failures have a higher success rate (57.1% fail-to-pass rate) compared to other groups. We did not see much impact of focal localization on the performance, with 31.9% and 29.1% for correct and incorrect localization, respectively. Test localization has a significant impact on the performance (36.9% vs. 18.2% for correct and incorrect localization). In our ablation study, we also found that test localization is more important than focal localization. In Appendix E, we discuss the impact of hallucination replacer in our pipeline.
Perspective | Category | # of Sample | #of fail-to-pass | fail-to-pass rate |
Type of Test | PatchExisting | 122 | 50 | 41.0 |
AdditionOnly | 327 | 91 | 27.8 | |
Test on old Codebase | Pass | 69 | 0 | 0 |
AssertionFail | 170 | 97 | 57.1 | |
Fail | 101 | 40 | 39.6 | |
Error | 109 | 4 | 3.7 | |
Focal Localization | Correct | 370 | 118 | 31.9 |
Wrong | 79 | 23 | 29.1 | |
Test Localization | Correct | 317 | 117 | 36.9 |
Wrong | 132 | 24 | 18.2 |
Heterogenous prompts vs. temperature.
Otter++ scaled well with samples up to 5, giving 0.5%-5.6% improvement (see Appendix C). Otter++ uses heterogeneous prompting instead of higher temperature to generate multiple samples. We make multiple LLM calls in different stages and multi-sampling in each stage would exponentially increase the test counts. Therefore, to compare heterogenous prompting with temperature, we generated 5 samples at high temperature (1.0) in the last LLM call in the Test Generator phase of Otter (our best solution). Though the average number of fail-to-pass test goes up (117.8 vs. 116.6), the Top-5 and Top-1 (using Otter++’s ranker) results remain lower (Top-5: 173 vs. 197 and Top-1: 146 vs. 166). That means heterogenous prompting boosts up the overall ensemble performance if we compare at the same number of samples. Appendix D discusses more details.
@@ 6.7 Test Generation and SWE Agents
The tests generated by our approach can be used for filtering bad code patches and increasing the precision of solutions proposed by different systems from the SWE-bench Verified leaderboard. We take the top 22 systems from the leaderboard and run the five tests generated by Otter++. We filter out a code patch if all the tests fail on it. Figure 6 shows that this achieves a precision of 65% to 92% while maintaining a decent recall of 30%-41%, except for one system where the precision increased by 22% to 167%. Note that Mündler et al. (2024) achieved 47.8% precision at 20% recall on SWE-Agent. Using tests generated by Otter++ achieved much higher precision while maintaining greater recall.
Apart from filtering SWE-patches, we could use our tests to choose the best SWE-patch, which would be a good application of the Otter-generated tests. We have tried CodeT ranking on candidates from the top 3 leaderboard systems on SWE-Bench-Verified and observed 2% improvement. Note that some of these leaderboard solutions have already been through good rankers and used superior models. Improving upon these samples using a ranker may be difficult.
@@ 6.8 Effect of Data Contamination
We conducted another experiment following the approach of Schäfer et al. (2024). We compute the similarity score between the generated test from Otter and the most similar test from the repository, as follows: , where is the set of test functions and is the generated test. Figure 8 shows the similarity scores for new tests. For 90% of the instances, the similarity score is less than 0.6. Table 12 in the appendix shows some samples with more than 0.5 similarity to give the reader some idea. From our observation, even at similarity score of 0.7, the tests are significantly different. As expected, for modified tests, the similarity is higher. However, we did not observe any difference between fail-to-pass tests and other tests (see appendix for figures). Therefore, the model is not simply generating memorized tests.
@@ 8 Conclusion
The primary contribution of this paper is Otter, a system for generating tests from issue descriptions before issue resolution. Otter outperforms the prior state of the art in fail-to-pass tests generated while also costing less. This paper also contributes TDD-Bench-Verified, a new benchmark for the same problem statement, mined from real-world GitHub issues with strict filters and evaluation metrics. Finally, this paper demonstrates that generated tests can improve patches generated by SWE agents, helping them reach a precision of between 65% and 92%. TDD-Bench-Verified and Otter generated tests are at https://github.com/IBM/TDD-Bench-Verified.
@@
This paper presents work whose goal is to advance the field of software testing and program repair. Developers spend a significant amount of time resolving bugs and testing them. We believe this work will significantly improve the developers’ experience in their day-to-day life. We envision two ways of integrating Otter into existing developer workflows. First, Otter can run on a new issue to propose a test, which the stakeholders can use to clarify requirements for the desired behavior after issue resolution and, following that, a developer can use the test for test-driven development to resolve the issue. Second, Otter can be paired with a patch-generation solution (a “SWE agent”) to create a PR that includes both a patch and a test.
@@ Appendix A Different attributes of the TDD-Bench-Verified Instances & Sample Test Patch
*File counts are based on the main branches of the project (cloned on October 29, 2024). |
@@ Appendix B Action Counts in Self-Reflective Action Planner Phase
Appendix C Scaling of Otter++
Otter++ uses heterogeneous prompting and execution logs to select the best solution. Therefore, we cannot significantly increase the number of samples. Otter++ scaled well with samples up to 5, giving 0.5%-5.6% improvement (see Table 9).
# of Candidate | # of Fail-to-pass | In % |
1 | 141 | 31.4 |
2 | 154 | 34.3 |
3 | 157 | 35.0 |
4 | 164 | 36.5 |
5 | 166 | 37.0 |
Appendix D Heterogenous Prompts vs Temperature
In Table 10, the Top-1 column shows “fail-to-pass @ 1” based on our actual ranker, whereas the Oracle column shows “fail-to-pass @ N”, i.e., the result if we had a perfect ranker. Heterogeneous prompts yield better results than high-temperature samples in both the Top-1 column and the Oracle column. This is evidence that the difference indeed comes from the diversity of heterogeneous prompts and is not ranker specific. The difference between the Top-1 and the Oracle column is almost the same in the first three rows (i.e., all settings with 5 samples). That said, there is still room for improvement in the ranker, as illustrated by the last row. In the last row, an oracle that could reliably choose among all 10 samples would perform amazingly well, motivating further work on better rankers.
Setting | Individual Performance | Max | Average | Top-1 | Oracle |
5 heterogeneous prompts (Otter++) | 141, 110,115, 107, 110 | 141 | 116.6 | 166 | 197 |
5 high-temperature samples (temp=1.0) | 111, 118, 126, 121, 113 | 126 | 117.8 | 146 | 173 |
5 = 1 greedy + 4 high-temperature samples | 141, 111, 118, 126, 121 | 141 | 123.4 | 152 | 183 |
10 = 5 heterogeneous + 5 high-temperature | 141, 110,115, 107, 110, 111, 118, 126, 121, 113 | 141 | 117.2 | 168 | 218 |
Appendix E Impact of hallucination replacer.
The localizer makes two LLM calls during which hallucinations can occur. In the first call, we drop hallucinated file names. In the second call, we replace hallucinated file names with existing ones to ensure our pipeline functions correctly. Table 11 show that our proposed technique helps Otter achieve better performance. For example, if we didn’t replace the hallucinated file name in the 2nd call of test localization (the last row), we would have lost 5 samples because our pipeline would have exited with an error. Replacing the hallucinated name generated 4 fail-to-pass tests for those samples.
Focal/Test | LLM Call |
#of fail-to-pass |
Focal | LLM Call 1 | 19 | 8 | 42.1 | ||||
LLM Call 2 | 5 | 3 | 60.0 | |||||
Test | LLM Call 1 | 39 | 22 | 56.4 | ||||
LLM Call 2 | 5 | 4 | 80.0 |
Appendix F Data Contamination
@@
Appendix G Prompts for Otter